# Adam Exploits $\ell_{\infty}$-geometry of Loss Landscape via Coordinate-wise Adaptivity

### v1

@@ 1 Introduction
Large language models (LLMs) have gained phenomenal capabilities as their scale grows (Radford et al., 2019; Kaplan et al., 2020; Brown et al., 2020; Zhang et al., 2022a; Touvron et al., 2023; OpenAI, 2023; Reid et al., 2024). However, pre-training LLMs are incredibly time-consuming. Adaptive Momentum Estimation ($\mathtt{Adam}$)(Kingma & Ba, 2014) is the current to-go optimization algorithm for LLMs due to its fast convergence. In contrast, $\mathtt{SGD}$, a popular and arguably the simplest optimizer, optimizes language model loss much slower than $\mathtt{Adam}$.
@@
However, the optimization benefit of $\mathtt{Adam}$ over $\mathtt{SGD}$ cannot be explained by existing theory. Existing convergence analyses for $\mathtt{Adam}$ and $\mathtt{SGD}$ focus on the dependence on the number of steps under assumptions on the smoothness and gradient bounds of the loss (Défossez et al., 2022), and it has been shown that both $\mathtt{Adam}$ and $\mathtt{SGD}$ achieve the minimax convergence rate $\widetilde{O}(T^{-1/4})$ in the non-convex settings (Arjevani et al., 2023). Thus according to the theory, in the worst case, $\mathtt{SGD}$ would be more desirable compared to $\mathtt{Adam}$ because they have the same convergence rate, and yet $\mathtt{Adam}$ is less memory-efficient due to its coordinate-wise adaptivity, which needs to store the empirical moving average of second-order moments of past stochastic gradients. Therefore, we hypothesize that the coordinate-wise adaptivity in $\mathtt{Adam}$ is exploiting some unknown properties of LLMs which $\mathtt{SGD}$ cannot make use of.
@@
Towards this end, we identified a significant difference between $\mathtt{Adam}$ and $\mathtt{SGD}$ in this paper. This difference, often ignored in previous works, is that $\mathtt{SGD}$ is rotation-invariant, while $\mathtt{Adam}$ is only permutation-invariant (see definitions in Section 2). Intuitively, this means if we rotate the loss landscape, the optimization trajectory of $\mathtt{SGD}$ would be the same (up to some rotation), while the trajectory of $\mathtt{Adam}$ could be completely different. If $\mathtt{Adam}$ optimizes much slower more slowly after rotation, then it suggests $\mathtt{Adam}$ is exploiting some non-rotational-invariant properties of the loss function, which is not captured by standard smoothness assumptions in the convergence analysis.
@@
Figure 1 summarizes our findings by comparing $\mathtt{Adam}$ on the original and rotated loss. The performance of $\mathtt{Adam}$ on rotated loss does become much worse than $\mathtt{Adam}$ on the original loss. We also test a memory-efficient and rotational-invariant variant of $\mathtt{SGD}$, $\mathtt{AdaSGD}$ (Wang & Wiens, 2020) (defined in Algorithm 2) [^1] There is one small difference. We use an exponential average of the gradient for ${\bm{m}}_{t}$ instead of momentum. Our definition makes $\mathtt{AdaSGD}$ the same as $\mathtt{Adam}$ in a one-dimensional problem.. Surprisingly, the rotated $\mathtt{Adam}$ performs even much worse than the $\mathtt{SGD}$ variant. The results suggest it is impossible to explain the superior optimization performance of Adam over SGD just using rotationally invariant assumptions on the loss function, which raises the natural question,
@@
We hypothesize that the $\ell_{2}$ lipschitzness of loss gradient does not provide a tight-enough characterization of loss landscape of deep learning models in practice, such that we can separate $\mathtt{Adam}$ and other rotational-invariant algorithms. Inspired by the similarity between $\mathtt{Adam}$ and $\mathtt{SignGD}$ and the fact that $\mathtt{SignGD}$ is the normalized steepest descent with respect to $\ell_{\infty}$ norm, we propose to use $\ell_{\infty}$-norm related smoothness as a better tool to analyze $\mathtt{Adam}$. In particular, our main results use the $(1,1)$-norm of Hessian of loss divided by variable dimension $d$ in replacement of the spectral norm of Hessian as the smoothness measure, and prove a convergence rate of $O(\frac{1}{\sqrt{T}})$ for $\mathtt{Adam}$ without noise, or $O((\frac{\log T}{T})^{1/4})$ with noise. Our results have the same dependence on $T$ as previous results, but much smaller smoothness constant when we measure it empirically. We empirically verify that $(1,1)$-norm of Hessian positively correlates with final training loss of $\mathtt{Adam}$ on both synthetic tasks like quadratic loss and real tasks like training GPT2 on OpenWebText and ResNet on CIFAR10.
@@
    We show by experiments that the empirical optimization advantage of $\mathtt{Adam}$ over $\mathtt{SGD}$ can not explained solely under rotation-invariant assumptions. (Figure 1)
@@
    We further generalize the theoretical analysis for $\mathtt{Adam}$ to blockwise $\mathtt{Adam}$ (Algorithm 3) whose convergence rate can be characterized by a novel smoothness measure (Theorem 3.11). $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ are two notable examples of blockwise $\mathtt{Adam}$. In $\mathtt{Adam}$, all blocks are of size $1$. In $\mathtt{AdaSGD}$, there is only one block.
@@
initialization ${\bm{x}}_{0}$, stochastic loss functions $\{L_{t}\}_{t=1}^{T}$
@@
initialization ${\bm{x}}_{0}$, stochastic loss functions $\{L_{t}\}_{t=1}^{T}$
@@
   $m_{t,i}\leftarrow\beta_{1}m_{t-1,i}+(1-\beta_{1})g_{t,i}$
@@
2 Preliminaries
Notations. For ${\bm{x}}\in\mathbb{R}^{d}$, we define the vector $p$-norm $\|{\bm{x}}\|_{p}$ as $(\sum_{i=1^{d}}x_{i}^{p})^{1/p}$ for $p\in[1,\infty]$. For a matrix $\bm{A}\in\mathbb{R}^{d_{1}\times d_{2}}$, its $(1,1)$-norm $\|\bm{A}\|_{1,1}$ is defined as $\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\left|A_{i,j}\right|$ and its operator norm induced by vector $p$-norm $\|\cdot\|_{p}$ as $\sup_{{\bm{x}}\in\mathbb{R}^{d}}\frac{\|\bm{A}{\bm{x}}\|_{q}}{\|{\bm{x}}\|_{p}}$, denoted by $\|\bm{A}\|_{p}$, where $\frac{1}{q}+\frac{1}{p}=1$ and $\|\cdot\|_{q}$ is the dual norm of $\|\cdot\|_{p}$. For a deterministic loss function $L({\bm{x}})$, we consider optimization over $L$ with only access independent stochastic functions $\{L_{t}({\bm{x}})\}_{t=1}^{T}$ such that $\mathbb{E}L_{t}({\bm{x}})=L({\bm{x}})$ for any input ${\bm{x}}\in\mathbb{R}^{d}$.
Rotation. For an invertible function $\mathcal{T}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, $\mathcal{T}$ is a rotating transformation if there exists an orthogonal matrix $\bm{T}\in\mathbb{R}^{d\times d}$ such that $\mathcal{T}({\bm{x}})=\bm{T}{\bm{x}}$. $\mathcal{T}$ is a permutating transformation if there exists a permutation $\pi:[d]\rightarrow[d]$ such that $\mathcal{T}({\bm{x}})=[x_{\pi(1)},\dots,x_{\pi(d)}]^{\top}$. A permutating transformation is always a rotating transformation. We will use $\mathcal{R}$ to denote a rotating transformation.
@@ 2 Preliminaries
An algorithm $A$ is invariant w.r.t. $\mathcal{T}$ if it always holds that $\tilde{{\bm{x}}}_{t}=\mathcal{T}^{-1}({\bm{x}}_{t})$ for any hyperparameters, initialization and stochastic losses. An algorithm $A$ is rotation invariant if it is invariant w.r.t. any rotating transformation $\mathcal{R}$. And $A$ is permutation invariant if it is invariant w.r.t. any permutating transformation.
The following Theorem 2.2 shows the difference between $\mathtt{Adam}$ and $\mathtt{AdaSGD}$, whose proof is in Appendix B.
@@
$\mathtt{SGD}$ and $\mathtt{AdaSGD}$ are rotation-invariant. $\mathtt{Adam}$ and $\mathtt{SignGD}$ are only permutation-invariant.
@@ 3 Main Results: Convergence Rates of $\mathtt{Adam}$
In this section, we present our main theoretical results, starting with a convergence analysis of $\mathtt{Adam}$ for stochastic smooth loss with coordinate-wise gradient noise (Theorem 3.5). We allow non-convex losses and thus the convergence is measured by the $\ell_{1}$ norm of the gradient. For deterministic loss, our best convergence rate (Theorem 3.2) is achieved by $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$). For stochastic loss with bounded gradient noise variance, our best rate (Corollary 3.6) is achieved by $\mathtt{RMSProp}$ ($\mathtt{Adam}$ with $\beta_{1}=0$ and $\beta_{2}\in[0,1]$).
Then we extend our analysis of $\mathtt{Adam}$ to more general blockwise $\mathtt{Adam}$ (Theorem 3.11), which contains both $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ as special cases. We also comes up with novel smoothness measures (Definition 3.9) corresponding to the set of blocks used in blockwise $\mathtt{Adam}$.
Similar to previous works (Défossez et al., 2022), our analysis could be extended to the most general case of $\mathtt{Adam}$, where both $\beta_{1}$, $\beta_{2}$ are non-zero, but the rate becomes strictly worse than the $\mathtt{RMSProp}$ (the case of $\beta_{1}=0$), as there will be some extra polynomials of $\frac{1}{1-\beta_{1}}$. We decide not to include result for the most general case, on one hand for ease of presentation, and on the other hand, because such result could explain the optimization benefit of momentum ($\beta_{1}>0$) in practice and does not add any insight on the benefit of $\mathtt{Adam}$. We hypothesis that we are missing some important features of loss landscape of transformers in the theoretical assumptions and we leave this for future work.
3.1 Warmup: $\mathtt{SignGD}$ ($\beta_{1}=\beta_{2}=0$)
In this section, we use the convergence analysis for $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$) as a warm-up and illustrate how $\mathtt{Adam}$ could benefit from a non-rotational invariant property of the loss landscape, which in particular is the $\ell_{\infty}$ smoothness. The key observation here is that $\mathtt{SignGD}$ is the normalized steepest descent with respect to $\ell_{\infty}$ norm (see (Xie & Li, 2024)), and thus it is more natural to analyze its convergence using $\ell_{\infty}$-norm-related geometry of the loss.
@@ 3.1 Warmup: $\mathtt{SignGD}$ ($\beta_{1}=\beta_{2}=0$)
Let $L$ be $H$-smooth with respect to $\ell_{\infty}$ norm and $\{{\bm{x}}_{t}\}_{t=1}^{T}$ be the iterates of $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$) on $L$ with initialization ${\bm{x}}_{0}$ and learning rate $\eta$, it holds that
$\min_{1\leq t\leq T}\left\|\nabla L({\bm{x}}_{t})\right\|_{1}\leq\frac{L({\bm{x}}_{0})-\min L}{T\eta}+\frac{H\eta}{2}$
if we choose $\eta=\sqrt{\frac{2(L({\bm{x}}_{0})-\min L)}{TH}}$, then $\min_{1\leq t\leq T}\left\|\nabla L({\bm{x}}_{t})\right\|_{1}\leq\sqrt{\frac{2H(L({\bm{x}}_{0})-\min L)}{T}}$.
3.2 Main result: $\mathtt{RMSProp}$ ($\beta_{1}=0,\beta_{2}\in[0,1]$)
It is well-known that $\mathtt{SignGD}$ might not converge in the stochastic case as the expectation of descent direction for mini-batch loss may not be a descent direction, and $\mathtt{RMSProp}$ is proposed to address this issue by using a moving average of the squared gradient per coordinate to reduce the coorleation between the denominator and the numerator, thus making the expected update direction less biased (Hinton et al., 2012). In this subsection we formalize the above intuition and show indeed a positive $\beta_{2}$ in $\mathtt{Adam}$ helps convergence in the stochastic case. The main challenges here are from both lower bounding the first-order term and upper bounding the second-order term in the modified descent lemma (the counterpart of Appendix A for $\mathtt{RMSProp}$).

### v3

@@ 1 Introduction
Large language models (LLMs) have gained phenomenal capabilities as their scale grows (Radford et al., 2019; Kaplan et al., 2020; Brown et al., 2020; Zhang et al., 2022a; Touvron et al., 2023; OpenAI, 2023; Reid et al., 2024). However, pre-training LLMs is incredibly time-consuming. $\mathtt{Adam}$ (Kingma & Ba, 2014) is the current to-go optimization algorithm for LLMs due to its fast convergence. In contrast, $\mathtt{SGD}$, a popular and arguably the simplest optimizer, optimizes language model loss much more slowly than $\mathtt{Adam}$.
@@
However, the optimization benefit of $\mathtt{Adam}$ over $\mathtt{SGD}$ cannot be explained by existing theory. Current convergence analyses for $\mathtt{Adam}$ and $\mathtt{SGD}$ focus on the dependence on the number of steps under assumptions on the smoothness of the loss and gradient bounds (Défossez et al., 2022), and it has been shown that both $\mathtt{Adam}$ and $\mathtt{SGD}$ achieve the minimax convergence rate $\widetilde{O}(T^{-1/4})$ in the non-convex settings (Arjevani et al., 2023). Thus according to the theory, in the worst case, $\mathtt{SGD}$ would be more desirable compared to $\mathtt{Adam}$ because they have the same convergence rate, and yet $\mathtt{Adam}$ is less memory-efficient due to its coordinate-wise adaptivity, which needs to store the empirical moving average of second-order moments of past stochastic gradients. Therefore, we hypothesize that the coordinate-wise adaptivity in $\mathtt{Adam}$ is exploiting some unknown properties of LLMs which $\mathtt{SGD}$ cannot make use of.
@@
Towards this end, we identified a big difference between $\mathtt{Adam}$ and $\mathtt{SGD}$, which is ignored in previous works. That is, $\mathtt{SGD}$ is rotation-equivariant, while $\mathtt{Adam}$ is only permutation equivariant (Definition 2.1). Intuitively, if we rotate the loss landscape, the trajectory of $\mathtt{SGD}$ would be the same (up to some rotation), while the trajectory of $\mathtt{Adam}$ can be completely different. If $\mathtt{Adam}$ optimizes much more slowly after rotation, it suggests $\mathtt{Adam}$ is exploiting some non-rotation-invariant properties of the loss, which is not captured by standard smoothness assumptions in the convergence analysis.
@@
Figure 1 summarizes our findings by comparing $\mathtt{Adam}$ on the original and rotated loss. The performance of $\mathtt{Adam}$ on the rotated loss does become much worse than $\mathtt{Adam}$ on the original loss. We also test a memory-efficient and rotation-equivariant variant of $\mathtt{SGD}$, $\mathtt{AdaSGD}$ (Wang & Wiens, 2020), defined in Algorithm 2. Surprisingly, the rotated $\mathtt{Adam}$ performs even much worse than the $\mathtt{SGD}$ variant. The results suggest it is impossible to explain the superior optimization performance of Adam over SGD just using rotation-invariant assumptions on the loss function, which raises the natural question,
@@
We hypothesize that the $\ell_{2}$-lipschitzness of loss gradient does not provide a tight-enough characterization of loss landscape of deep learning models in practice, such that we can separate $\mathtt{Adam}$ and other rotation-equivariant algorithms. Inspired by the similarity between $\mathtt{Adam}$ and $\mathtt{SignGD}$ and the fact that $\mathtt{SignGD}$ is the normalized steepest descent with respect to $\ell_{\infty}$-norm, we propose to use $\ell_{\infty}$-norm related smoothness as a better tool to analyze $\mathtt{Adam}$. In particular, our main results use the $(1,1)$-norm of the Hessian of the loss normalized by variable dimension $d$, as the smoothness measure, instead of its spectral norm. And we prove a convergence rate of $O(\frac{1}{\sqrt{T}})$ for $\mathtt{Adam}$ without noise, or $O((\frac{\log T}{T})^{1/4})$ with noise. Our results have the same dependence on $T$ as previous results, but a much smaller smoothness constant when measured empirically. We empirically verify $(1,1)$-norm of Hessian positively correlates with final training loss of $\mathtt{Adam}$ on both synthetic tasks like quadratic loss and real tasks like training GPT2 on OpenWebText and ResNet on CIFAR10.
@@
    We show by experiments that the empirical optimization advantage of $\mathtt{Adam}$ over $\mathtt{SGD}$ can not be explained solely under rotation-invariant assumptions. (Figure 1)
@@
    We further generalize the theoretical analysis  (Theorem 3.11) for $\mathtt{Adam}$ to blockwise $\mathtt{Adam}$ (Algorithm 3) whose convergence rate can be characterized by a novel smoothness measure (Definition 3.9). $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ are two notable examples of blockwise $\mathtt{Adam}$. In $\mathtt{Adam}$, all blocks are of size $1$. In $\mathtt{AdaSGD}$, there is only one block.
@@
2 Preliminaries
Notations.
For ${\bm{x}}\in\mathbb{R}^{d}$, we define the vector $p$-norm $\|{\bm{x}}\|_{p}$ as $(\sum_{i=1}^{d}x_{i}^{p})^{1/p}$ for $p\in[1,\infty]$. For a matrix $\bm{A}\in\mathbb{R}^{d_{1}\times d_{2}}$, its $(1,1)$-norm $\|\bm{A}\|_{1,1}$ is defined as $\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\left|A_{i,j}\right|$ and its operator norm induced by vector $p$-norm $\|\cdot\|_{p}$ as $\sup_{{\bm{x}}\in\mathbb{R}^{d}}\frac{\|\bm{A}{\bm{x}}\|_{q}}{\|{\bm{x}}\|_{p}}$, denoted by $\|\bm{A}\|_{p}$, where $\frac{1}{q}+\frac{1}{p}=1$ and $\|\cdot\|_{q}$ is the dual norm of $\|\cdot\|_{p}$. For a square matrix $\bm{A}\in\mathbb{R}^{d\times d}$, $\left|\bm{A}\right|$ is defined as the unique square root of $\bm{A}^{\top}\bm{A}$. For a deterministic loss function $L({\bm{x}})$, we consider optimization over $L$ with access only to independent stochastic functions $\{L_{t}({\bm{x}})\}_{t=1}^{T}$ such that $\mathbb{E}L_{t}({\bm{x}})=L({\bm{x}})$ for any input ${\bm{x}}\in\mathbb{R}^{d}$.
Rotation.
For an invertible function $\mathcal{T}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, $\mathcal{T}$ is a rotating transformation if there exists an orthogonal matrix $\bm{T}\in\mathbb{R}^{d\times d}$ such that $\mathcal{T}({\bm{x}})=\bm{T}{\bm{x}}$. $\mathcal{T}$ is a permutating transformation if there exists a permutation $\pi:[d]\rightarrow[d]$ such that $\mathcal{T}({\bm{x}})=[x_{\pi(1)},\dots,x_{\pi(d)}]^{\top}$. A permutating transformation is always a rotating transformation. We will use $\mathcal{R}$ to denote a rotating transformation.
@@ 2 Preliminaries
initialization ${\bm{x}}_{0}$, stochastic losses $\{L_{t}\}_{t=1}^{T}$
@@
initialization ${\bm{x}}_{0}$, stochastic losses $\{L_{t}\}_{t=1}^{T}$
@@
   $m_{t,i}\leftarrow\beta_{1}m_{t-1,i}+(1-\beta_{1})g_{t,i}$ [^2] Here it is slightly different from Wang & Wiens (2020). We use an exponential average of the gradient for ${\bm{m}}_{t}$ instead of momentum. Our definition makes $\mathtt{AdaSGD}$ the same as $\mathtt{Adam}$ in a one-dimensional problem.
@@
An algorithm $A$ is equivariant w.r.t. $\mathcal{T}$ if it always holds that $\tilde{{\bm{x}}}_{t}=\mathcal{T}^{-1}({\bm{x}}_{t})$ for any hyperparameters, initialization and stochastic losses. An algorithm $A$ is rotation-equivariant if it is equivariant w.r.t. any rotating transformation $\mathcal{R}$. And $A$ is permutation-equivariant if it is equivariant w.r.t. any permutating transformation.
The following Theorem 2.2 shows the difference between $\mathtt{Adam}$ and $\mathtt{AdaSGD}$, whose proof is in Appendix A. We provide a visual example in Figure 2. It also shows the similarity between $\mathtt{SignGD}$ and $\mathtt{Adam}$.
Figure 2: Optimization trajectories of different algorithms on a quadratic function before (left) and after (right) rotating the coordinate system. The starting point is marked in black. $\mathtt{AdaSGD}$ and $\mathtt{SGD}$ exhibit rotation-equivariant behavior, as their trajectories rotate consistently with the objective function. In contrast, $\mathtt{Adam}$ and $\mathtt{SignGD}$ fail to preserve trajectory shape under rotation, demonstrating that they are not rotation-equivariant.
@@
$\mathtt{SGD}$ and $\mathtt{AdaSGD}$ are rotation-equivariant. $\mathtt{Adam}$ and $\mathtt{SignGD}$ are only permutation-equivariant.
@@ 3 Main Results: Convergence Rates of $\mathtt{Adam}$
In this section, we present our main theoretical results, starting with a convergence analysis of $\mathtt{Adam}$ for stochastic smooth losses with coordinate-wise gradient noise (Theorem 3.5). We allow non-convex losses and thus the convergence is measured by the $\ell_{1}$ norm of the gradient. For the deterministic loss, our best convergence rate (Theorem 3.2) is achieved by $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$). For the stochastic loss with bounded gradient noise variance, our best rate (Corollary 3.6) is achieved by $\mathtt{RMSProp}$ ($\mathtt{Adam}$ with $\beta_{1}=0$ and $\beta_{2}\in[0,1]$).
Then we extend our analysis of $\mathtt{Adam}$ to more general blockwise $\mathtt{Adam}$ (Theorem 3.11), which contains both $\mathtt{Adam}$ and $\mathtt{AdaSGD}$ as special cases. We also come up with novel smoothness measures (Definition 3.12) corresponding to the set of blocks used in blockwise $\mathtt{Adam}$.
Similar to previous works (Défossez et al., 2022), our analysis could be extended to the most general case of $\mathtt{Adam}$, where both $\beta_{1}$, $\beta_{2}$ are non-zero. But the rate becomes strictly worse than $\mathtt{RMSProp}$ (the case of $\beta_{1}=0$), as there will be some extra polynomials of $\frac{1}{1-\beta_{1}}$. We decide not to include the result for the most general case, on the one hand for ease of presentation, and on the other hand, because such result can not explain the optimization benefit of momentum ($\beta_{1}>0$) in practice and does not add any insight on the benefit of $\mathtt{Adam}$. We hypothesize that we are missing some important features of loss landscape of transformers in the theoretical assumptions and we leave this for future work.
3.1 Warmup: $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$)
In this section, we present the convergence analysis for $\mathtt{SignGD}$ as a warm-up and illustrate how $\mathtt{Adam}$ could benefit from a non-rotation-invariant property of the loss landscape, which in particular is the $\ell_{\infty}$ smoothness. The key observation here is that $\mathtt{SignGD}$ is the normalized steepest descent with respect to $\ell_{\infty}$ norm (Xie & Li, 2024), so it is more natural to analyze its convergence using $\ell_{\infty}$ norm geometry of the loss.
@@ 3.1 Warmup: $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$)
Theorem 3.2 gives the convergence rate for $\mathtt{SignGD}$ and the proof is in Appendix B.
@@
Let $L$ be $H$-smooth w.r.t. $\ell_{\infty}$ norm and $\{{\bm{x}}_{t}\}_{t=1}^{T}$ be the iterates of $\mathtt{SignGD}$ ($\mathtt{Adam}$ with $\beta_{1}=\beta_{2}=0$) on $L$ with initialization ${\bm{x}}_{0}$ and learning rate $\eta$, it holds that
$\min_{1\leq t\leq T}\left\|\nabla L({\bm{x}}_{t})\right\|_{1}\leq\frac{L({\bm{x}}_{0})-\min L({\bm{x}})}{T\eta}+\frac{H\eta}{2}.$
If we choose $\eta=\sqrt{\frac{2(L({\bm{x}}_{0})-\min L({\bm{x}}))}{TH}}$, then $\min_{1\leq t\leq T}\left\|\nabla L({\bm{x}}_{t})\right\|_{1}\leq\sqrt{\frac{2H(L({\bm{x}}_{0})-\min L({\bm{x}}))}{T}}$.
3.2 Main result: $\mathtt{RMSProp}$ ($\mathtt{Adam}$ with $\beta_{1}=0,\beta_{2}\in[0,1]$)
It is well-known that $\mathtt{SignGD}$ might not converge in the stochastic case as the expectation of descent direction for mini-batch loss may not be a descent direction for $L$. $\mathtt{RMSProp}$ is proposed to address this issue by using a moving average of the squared gradient per coordinate to reduce the correlation between the denominator and the numerator, thus making the expected update direction less biased (Hinton et al., 2012). In this subsection we formalize the above intuition and show indeed a positive $\beta_{2}$ in $\mathtt{Adam}$ helps convergence in the stochastic case. The main challenges here are from both lower bounding the first-order term and upper bounding the second-order term in the modified descent lemma (the $\mathtt{RMSProp}$ counterpart of Appendix B).